OpenStudy (anonymous):

Does 2sinxtanx = 3 === 2-2cos^2x over 2cosx=3?

OpenStudy (anonymous):

\[{2-2\cos^2x \over 2cosx}=3\]

sam (.sam.):

You want to find the solution? or just proving?

OpenStudy (anonymous):

prove and find the solution. But is 2sinxtanx equivalent to what I wrote?

OpenStudy (anonymous):

the answer to your question is if they're equivalent is no. cross out the leading 2 and it will be.

OpenStudy (anonymous):

but that's not the route i'd take...

OpenStudy (anonymous):

all yours sam..

sam (.sam.):

\[\frac{2(1-\cos^{2}x)}{2cosx}=3\] \[\frac{(1-\cos^{2}x)}{cosx}=3\] \[1-\cos^{2}x=3cosx\] \[-\cos^{2}x-3cosx+1=0\]

sam (.sam.):

No whole number solutions

OpenStudy (anonymous):

Hmm Ok... But there should be... I think

sam (.sam.):

Solution for that is \[x=\sin ^{-1}\left(\sqrt{\frac{3}{2} \left(\sqrt{13}-3\right)}\right)\approx 1.26319\] and \[x=-\sin ^{-1}\left(\sqrt{\frac{3}{2} \left(\sqrt{13}-3\right)}\right)\approx -1.26319\]

OpenStudy (anonymous):

Won't it be 2sinxtanx = 3 => 2(sin^2x)/cosx = 3 => 2(1 - cos^2x)/cosx = 3 => 2 - 2cos^2x = 3cosx => 2cos^2x + 3cosx - 2 = 0 => (2cosx - 1)(cosx + 2) = 0 cosx = 1/2, cosx = -2 => Impossible => x = Pi/3, 2Pi - Pi/3 = 5Pi/3

sam (.sam.):

The "2" canceled from start , the furthest you can get is sin(x) tan(x)=3

OpenStudy (callisto):

2sinxtanx is not the same as 2-2cos^x over 2cosx

OpenStudy (anonymous):

Ok. Is it the same as 2-2cos^2x over cosx?

OpenStudy (callisto):

yes

OpenStudy (anonymous):

Thank you :)

OpenStudy (callisto):

You're welcome.. I didn't help anyway...

OpenStudy (anonymous):

@c How do you get from the ranges 0<x<360 with the answer cosx=1/2? The first is 60, and then what?

OpenStudy (anonymous):

@Callisto

OpenStudy (callisto):

 Very ugly drawing sorry. But from the graph, we can know there are 2 solutions for cosx =0.5 as you know the first one is 60 the second solution is in quadrant IV , which gives positive value for cosx (i meant it is like the quadrant I) and from cos (360 -x) = cos x, you can get the 2nd angle by 360 -x =60 Hmm.. is that your question?

OpenStudy (anonymous):

Yes, thank you :)

OpenStudy (callisto):

you're welcome!